Johnson Products earned $5.20 per share last year and paid a $2.15 per share dividend. If ROE was 16 percent, what is the sustainable growth rate? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Omit the "%" sign in your response.)

   

  Sustainable growth rate %
Answer #1

Sustainable Growth Rate = ROE * (1- dividend payout ratio)

Dividend Payout Ratio = Dividend/ Earning Per Share

Dividend Payout Ratio = 2.15/5.2 = 41%

Sustainable Growth Rate = 16*(1-.41)

Sustainable Growth Rate = 9.44%